OpenStudy (anonymous):

help pleaseee?:/ The lengths of the diagonals of a rhombus are 2 in. and 5 in. find the measures of the angles of the rhombus to the nearest degree.

OpenStudy (anonymous):

idk sorry!

OpenStudy (anonymous):

ook thaaankss:)

OpenStudy (anonymous):

you are welcome

OpenStudy (whpalmer4):

Which class is this for? I'm a bit skeptical that they are asking you to solve this with ruler and protractor...

OpenStudy (whpalmer4):

Also, the instructions say the diagonals are 2" and 5", not the sides...

OpenStudy (anonymous):

geometry

OpenStudy (whpalmer4):

@kelly226 what are you getting for your answers? @lilkittykat36's answers are NOT correct. You do not even have to draw the figure to know this: the sum of the interior angles of a rhombus should add to 360 degrees.

OpenStudy (anonymous):

62 and 110 r the answers.. i already submitted it

OpenStudy (whpalmer4):

Not the correct answers!

OpenStudy (anonymous):

how?

OpenStudy (whpalmer4):

The diagonals of a rhombus intersect at 90 degree angles. They also bisect the angles. That means that for each of the 4 triangles that make up the rhombus, the angles are 90, 1/2 of one angle and 1/2 of the other angle. 90 + 1/2 (62) + 1/2(110) = 90 + 31 + 55. That does not add up to 180, but the sum of the interior angles of a triangle is always 180 degrees. The correct answers can be shown with a bit of trigonometry to be 43.6 degrees for the narrow angle and 136.4 degrees for the wider angle. Rounded to the nearest degree, 44 and 136. You'll note that 136+136+44+44 = 272+88 = 360, just as it should. Sum of the interior angles of a 4-sided polygon is always 360 degrees.